OpenStudy (anonymous):

Find two integers that make the equation true y2 =25 true

OpenStudy (accessdenied):

So, what two numbers when squared will equal 25? Remember that, if you square a negative number, it still becomes positive. ;)

OpenStudy (accessdenied):

Well, if we square \(5\), we'll get 25. But since squaring a negative number gets you to positive, we would also be able to square the negative of it, right? \((\neg 5)^2 = (\neg 1)^2 \times 5^2 = 1\times 25 = 25\)